# CanadaFireSat: Toward high-resolution wildfire forecasting with multiple modalities

### v1

@@ 4.2 Performance Analysis
We evaluate the performance of the two different architectures, CNN and ViT, across three different settings described in Section 3: [Uncaptioned image]satellite images only, [Uncaptioned image]environmental predictors only, and [Uncaptioned image]satellite and environmental data. The results are reported in Table 4.
@@
Table 4: Performance comparison of different model settings.
Across the three evaluation sets (last column of Table 4), both ResNet-50 and ViT-S trained on [Uncaptioned image]satellite and environmental data reach the highest performance, with $+15\%$ PRAUC and $+11\%$ F1 score for the CNN, compared to the FWI baseline. For both the CNN and ViT architectures, relying on multi-modal inputs shows an improvement over models trained on [Uncaptioned image]satellite images only or [Uncaptioned image]environmental predictors only. While individually [Uncaptioned image]satellite images only and [Uncaptioned image]environmental predictors only are already highly discriminative, the multi-modal setting [Uncaptioned image]satellite and environmental data remains the most accurate forecaster with an average gain of $+1.8\%$ in PRAUC and $+1.3\%$ in F1 score for CNN-based models and $+1.1\%$ in PRAUC and $+1.7\%$ in F1 score for the ViT. We further discuss the role of satellite image time series and environmental predictors in Section 5.1. On the original Test set, the better performance of all models (compared to the Val set) can be explained by fire patterns likely being more easily distinguishable due to the extreme fire season, a behavior also observed in the FWI baseline. Nonetheless, our best performing CNN model relying on [Uncaptioned image]satellite and environmental data still outperforms the FWI baseline, on the Test set by $+13.9\%$ (PRAUC) and by $+10\%$ (F1 score). The drop in performance of all models on the Test Hard set demonstrates the impact of the sampling strategy and the necessity of such an evaluation set to provide a lower bound of performance, as the models were not trained on such adversarial negative samples. When it comes to the comparison between ResNet-50 and ViT-S trained on [Uncaptioned image]satellite and environmental data, the former shows to perform best in terms of F1 score across all sets. However, differences remain small, and both architectures seem valid solutions for wildfire forecasting.
[…]
@@
We also study the results of ResNet-50 and ViT-S trained on [Uncaptioned image]satellite and environmental data across the most common land cover classes in CanadaFireSat. Our models struggle the most on the classes wetland and cropland. Indeed, fire patterns in these two land cover types differ from those observed in the majority of wildfires, affecting forest areas. In particular, peatland fires in Canada can occur under the ground in wet areas, or even under the snow layer. Such fires are difficult to observe through the predictors considered in the proposed CanadaFireSat, and would require ad-hoc modeling due to the specificities of such ecosystems. Low scores are also observed for cropland fires, which also present unique fire patterns, as the ignition is often human-induced and driven by a specific need for agricultural practices.
Figure 10: Refer to captionsatellite and environmental data models’ performance across the main land cover classes.
@@ 4.3 Deployment at Scale: Case Study
We chose a large wildfire that occurred in Québec on $2023/07/05$ as a case study, illustrated in Figure 11. The first row displays the RGB composite of the region of interest of size $14\>\text{km}\times 26\>\text{km}$ acquired by Sentinel-2 right before the wildfire starts on $2023/06/28$. The fire scar polygons from NBAC are shown on the second row. The third row shows the predictions of the CNN model trained in the multi-modal setting [Uncaptioned image]satellite and environmental data, on the positive samples overlapping with the considered ROI. The predictions follow the same pattern as the actual wildfire, despite slightly overestimating its extent, as it can be seen on both sides of the Sentinel-2 tile. This case study showcases the potential of CanadaFireSat to enable the deployment of models capable of monitoring large regions at high-resolution (i.e., Sentinel-2 spatial resolution).
@@
Figure 11: Row 1 Sentinel-2 tile from $2023/06/28$ of size $14\>\text{km}\times 26\>\text{km}$ before a large wildfire in Québec. Row 2 Fire polygons for the large wildfire on $2023/07/05$ over the same tile. Row 3 Binary model predictions (in red) over the $2.64\>\text{km}\times 2.64\>\text{km}$ center-cropped positive samples outlined in black.
@@ 5.1 High-resolution Wildfire Forecasting via Multi-modal Learning
As previously demonstrated in (Pelletier et al., 2023; Chowdhury and Hassan, 2015; Yang et al., 2021), multi-spectral multi-temporal satellite data can be a valuable predictor for wildfire forecasting. Indeed, several spectral indexes discriminative for wildfire forecasting can be extracted from Sentinel-2: normalized difference vegetation index (NDVI), normalized difference water index (NDWI), tasseled cap wetness, and channel histograms. The results reported in Table 4 demonstrate the potential of multi-spectral temporal satellite data for high-resolution wildfire forecasting (setting [Uncaptioned image]satellite images only). While hydrometeorological data (setting [Uncaptioned image]environmental predictors only) are commonly used in global and continental wildfire forecasting models, they can be complemented by satellite data to improve strongly both the spatial resolution and the accuracy of the prediction (setting [Uncaptioned image]satellite and environmental data).
@@ 5.2 Importance of Negative Sampling for Training and Evaluation
Numerous wildfire forecasting benchmarks require sampling the negative (non-fire) samples due to extreme imbalance and computational constraints. A common strategy is to focus on samples that burned once across the period studied (Prapas et al., 2022; Bakke et al., 2023; Prapas et al., 2023). In CanadaFireSat, we opt for a different strategy as we sample negative examples for each Canadian province uniformly across their yearly fire driver patterns (FWI values). By sampling the training, validation, and test sets in this way, we aim to train and evaluate our models on a subset representative of the conditions encountered in the entire Canada. Nonetheless, as the yearly fire patterns vary, the distribution of negatives w.r.t. the FWI changes over the years, in turn affecting models’ performance. This motivated the creation of a second test set to understand the impact of sampling on the models’ performance: Test Hard, sampled differently by aiming to replicate as close as possible relevant fire drivers from the positive set. We consider Test Hard as a lower bound for our model performance, since the main predictors for differentiating positive from negative patches are linked to ignition, a complex phenomenon difficult to model (Chen et al., 2021; Calef et al., 2008). Indeed, despite identical hydrometeorological conditions, a random ignition caused by humans or lightning is still necessary to cause a wildfire. Theoretically, wildfire forecasting could be framed as a positive unlabeled learning task (Bekker and Davis, 2020) to explicitly model the differences between fire conditions and ignition in wildfire forecasting. Examples of positive unlabeled tasks can be found in several environmental and ecological applications, like species distribution modeling (Zbinden et al., 2024) and phenology event detection (Fasihi et al., 2025). Fine-tuning our multi-modal models on data such as Test Hard or enhancing our set of predictors with proxies of ignition probability (e.g, proximity to human settlements, or lightning probability) are relevant directions for improving our models towards accounting for ignition probability.
@@ 5.3 Modeling Wildfires in the Boreal Ecosystem
As initially stated in Section 1, one of our main motivations is the rise of wildfires in the boreal ecosystem and the risks this represents for its local communities. To cover the areas of interest and to evaluate the broader impact of wildfires on global climate, we created our benchmark CanadaFireSat so that it covers the entirety of Canada, including all its agricultural lands, urban areas, and other ecosystems such as the temperate forest in British Columbia. To study the behavior of trained models on the boreal ecosystem, it is possible to constrain the analysis on the main land cover classes of the boreal ecosystem (needleleaf forest and wetlands). With our benchmark models trained on CanadaFireSat, we observe an important difference of performance between those two land cover classes: with Multi-Modal CNN and ViT performing respectively $+11.9\%$ and $+11.4\%$ better on needleleaf forest compare to wetlands in terms of weighted F1 score on Val and Test. Indeed, wetland wildfires are a unique phenomenon compared to forest wildfires as they depend much more on soil-related predictors and can burn underground for a long period. As a consequence, they are sometimes undetectable for optical remote sensing satellites. In particular, peatland wildfires that emit large amounts of CO2 and mercury (Fraser et al., 2018; Kohlenberg et al., 2018) are commonly studied independently from forest fires (Pelletier et al., 2023; Bali et al., 2021). Extending CanadaFireSat so that it includes data acquired from radar remote sensing satellites (for instance, Sentinel-1 images) could help to better model the surface soil conditions for wetlands and bridge the gap in performance across the boreal ecosystem.
@@ 5.4 Operationalization of the Model
Deploying models trained with CanadaFireSat over the entirety of Canada would require densely sampling the country with Sentinel-2 image time series, resulting in a huge amount of data to be processed. Indeed, the proposed dataset is aimed at modeling wildfire patterns at a moderate scale, but at high-resolution, and can be coupled with coarser resolution approaches (Prapas et al., 2022; Bali et al., 2021) to identify areas of interest and then apply our model to map such areas more precisely. Such coupling would allow wildfire management experts to target specific areas at risk for fine-grained wildfire forecasting or focus on areas that require more surveillance due to their proximity to local communities or due to their ecological and environmental interest. Learning models directly capable of multi-scale prediction is an interesting future research direction to deploy high-resolution wildfire forecasting at scale. Such a model would exploit hierarchical learning approaches developed in computer vision for semantic segmentation (Li et al., 2022; Atigh et al., 2022).
@@ 5.5 Limitations and Future Work
As mentioned in Section 5.2, the main limitation of methods trained on CanadaFireSat is the difficulty of modeling the ignition component in wildfires due to its inherent stochasticity. Weather data from ERA5 can provide information on the risk of lightning, nonetheless, explicitly adding lightning probability (Geng et al., 2019) as a predictor, as well as other proxies for human ignition, like the proximity to human settlement, could help the trained models to characterize ignition better.
@@
Multi-task learning (Zhang and Yang, 2021) can also be leveraged to develop a model capable of forecasting wildfires at multiple scales. One could leverage different forecasting heads at multiple resolutions: $10$ km, $1$ km, $100$ m. This could help alleviate memory size constraints when high-resolution forecasts are deemed unnecessary and help providing consistent predictions across scales. Indeed, hierarchical learning (Li et al., 2022) can also be linked to multi-scale wildfire forecasting, as one can build a hierarchical structure across labels at different resolutions.
@@
Moreover, one could investigate the potential of geolocation embeddings such as SatCLIP (Klemmer et al., 2023) or GeoCLIP (Vivanco Cepeda et al., 2023) to represent high-resolution non-dynamic satellite information combined with non-spatial, but temporal dynamics from Sentinel-2 (Pelletier et al., 2023), as a way to factorize spatial and temporal components in satellite data and limit memory consumption.
@@
Another line of improvement would be relying on pre-trained foundation models for Earth observation. For our Multi-Modal architectures, we leverage image encoders pre-trained on natural images such as ImageNet or via DINOv2, which are very different from multi-spectral satellite images. With the drastic increase in availability of Earth observation data, several models are being proposed to learn in an unsupervised way generalizable representations from this data (Cong et al., 2022; Jakubik et al., 2023; Hong et al., 2024; Astruc et al., 2024). Those models are mainly driven by the concept of masked auto-encoders (He et al., 2022) and contrastive learning (Chen et al., 2020; He et al., 2020). One could study the potential of those foundation models on high-resolution wildfire forecasting by fine-tuning them on CanadaFireSat.
@@
Finally, the increased complexity of models raises concerns regarding their interpretability and the possibility of understanding the role of the input variables in the final predictions. Several approaches exist to provide interpretations of black box wildfire forecasting models via feature attributions (Sundararajan et al., 2017; Selvaraju et al., 2017) or ranking (Lundberg and Lee, 2017), or even to directly build interpretable wildfire forecasting model architectures (Koh et al., 2020; Chen et al., 2019) via dense prediction architecture (Sacha et al., 2023; Porta et al., 2025a). However, those methods need adaptation to accommodate multi-modal (Ekim and Schmitt, 2023; Wang et al., 2023) or multi-temporal data (Turbé et al., 2023; Gee et al., 2019; Ghosal and Abbasi-Asl, 2021). They are also often not directly applicable to Earth observation data (Porta et al., 2025b) due to strong implicit bias for natural images (Chen et al., 2019). This gap remains unfulfilled, and future works should explore interpretable methods specifically tailored to Earth observations problem.

### v2

@@ 4.1 Performance Analysis
We evaluate the performance of the two different architectures (CNN and ViT) across three different settings described in Section 3: [Uncaptioned image]satellite images only, [Uncaptioned image]environmental predictors only, and [Uncaptioned image]satellite and environmental data. The results are reported in Table 4.
@@
Table 4: Performance comparison of different model settings. Bold indicates the best metric value for each dataset split and model type, and underline denotes the runner-up.
Across the three evaluation sets (last column of Table 4), both ResNet-50 and ViT-S trained on [Uncaptioned image]satellite and environmental data reach the highest performance, with $15\%$ PRAUC and $11\%$ F1 score for the CNN, compared to the FWI baseline. For both the CNN and ViT architectures, relying on multi-modal inputs shows, on average, an improvement over models trained on [Uncaptioned image]satellite images only or [Uncaptioned image]environmental predictors only. While individually [Uncaptioned image]satellite images only and [Uncaptioned image]environmental predictors only are already highly discriminative, the multi-modal setting [Uncaptioned image]satellite and environmental data remains the most accurate forecaster with an average gain of $1.8\%$ in PRAUC and $1.3\%$ in F1 score for CNN-based models and $1.1\%$ in PRAUC and $1.7\%$ in F1 score for the ViT. We further discuss the role of satellite image time series and environmental predictors in Section 5.1. On the Test set, the better performance for both F1 score and PRAUC of all models compared to the Val set can be explained by fire patterns likely being more easily distinguishable due to the extreme fire season, a behavior also observed in the FWI baseline. Nonetheless, our best performing CNN model relying on [Uncaptioned image]satellite and environmental data still outperforms the FWI baseline, on the Test set by $13.9\%$ (PRAUC) and by $10\%$ (F1 score). Further analysis of the model’s performance threshold analysis across the Val, Test, and Test Hard sets is shown in D. The drop in performance of all models on the Test Hard set demonstrates the impact of the sampling strategy and the necessity of such an evaluation set: Test Hard can be used to assess models’ lower bound performance and their ability to model the hidden phenomena behind ignition. When it comes to the comparison between ResNet-50 and ViT-S trained on [Uncaptioned image]satellite and environmental data, the former shows to perform best in terms of F1 score across all sets. However, differences remain small, and both architectures seem valid solutions for wildfire forecasting.
The performances of the different models in setting [Uncaptioned image]satellite and environmental data are studied in Figure 9 for increasing FWI values. We first focus on the False Positive Rate, defined as $\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}}$. As expected, we can observe in Figure 9(a) a positive correlation between the FPR and the FWI. Indeed, negative samples associated with a higher FWI show similar fire danger conditions to positive samples, and are thus much more difficult to discriminate, with ignition becoming the main triggering factor for samples with $\text{FWI}>20$. Then, we study in Figure 9(b) the variations of the weighted F1 score, defined as $\hat{\text{F1}}=\frac{\text{F1}-\text{F1}_{pos}}{1-\text{F1}_{pos}}$. This second index tells how good the model is compared to a naive predictor: $\text{F1}_{pos}$, assigning the class fire to all samples. We note a negative correlation between $\hat{\text{F1}}$ and the FWI: as the FWI increases, there is approximately no difference between our benchmark models and a naive predictor. Such behavior is not surprising as it is more likely for a sample associated to a high FWI to belong to the class fire, as confirmed by the increase in percentage of positive samples with higher FWI (from $13\%$ at $\text{FWI}\in[0,5]$ to $77\%$ at $\text{FWI}\in[20,30])$. We can conclude that the improved performance of our model with respect to the FWI baseline reported in Table 4 is due to a better prediction of wildfire occurrence at lower FWIs, as the task becomes trivial for $\text{FWI}>20$ due to data imbalance.
@@
We also study the results of ResNet-50 and ViT-S trained on [Uncaptioned image]satellite and environmental data across the most common land cover classes in CanadaFireSat. Our models struggle the most on the classes wetland and cropland. Indeed, fire patterns in these two land cover types differ from those observed in the majority of wildfires, which tend to affect forest areas. In particular, peatland fires in Canada can occur under the ground in wet areas, or even under the snow layer. Such fires are difficult to observe through the predictors considered in the proposed CanadaFireSat, and would require ad-hoc modeling due to the specificities of such ecosystems. Low scores are also observed for cropland fires, which also present unique fire patterns, as the ignition is often human-induced and driven by a specific need for agricultural practices. As before, detecting these events seems hardly possible with our remote sensing-based system.
Figure 10: Refer to captionsatellite and environmental data models’ F1 score across the main land cover classes.
@@ 4.2 Deployment at Scale: Case Study
We chose as a case study a large wildfire that occurred in British Columbia on $2023/07/01$, illustrated in Figure 11. The first row displays the RGB composite of the region of interest of size $16\>\text{km}\times 22\>\text{km}$ acquired by Sentinel-2 right before the wildfire starts on $2023/06/06$. The fire scar polygons from NBAC are shown on the second row. The third row shows the binarized predictions of the CNN model trained in the multi-modal setting [Uncaptioned image]satellite and environmental data, on the positive samples overlapping with the considered ROI. We observe how well the model delineates the urban interface on the left side of the wildfire and the rough approximation of its boundaries on the right side of the fire. However, we can also see at the top of the Sentinel-2 image that the model overestimates the wildfire extent. This case study showcases the potential of CanadaFireSat to enable the deployment of models capable of monitoring large regions at the unprecedented resolution of $100$ m.
@@
Figure 11: Row 1 Sentinel-2 tile from $2023/06/06$ of size $16\>\text{km}\times 22\>\text{km}$ before a large wildfire in British Columbia. Row 2 Fire polygons for the large wildfire on $2023/07/01$ over the same tile. Row 3 Binary model predictions (in red) over the $2.64\>\text{km}\times 2.64\>\text{km}$ center-cropped positive samples (patches outlined in black).
@@ 5.1 High-resolution Wildfire Forecasting via Multi-modal Learning
As previously demonstrated in (Pelletier et al., 2023; Chowdhury and Hassan, 2015; Yang et al., 2021), multi-spectral multi-temporal satellite data can be a valuable data source to forecast wildfires. Indeed, several spectral indices discriminative for wildfire forecasting can be extracted from Sentinel-2: normalized difference vegetation index (NDVI), normalized difference water index (NDWI), tasseled cap wetness, and channel histograms. The results reported in Table 4 demonstrate the potential of multi-spectral temporal satellite data for high-resolution wildfire forecasting (setting [Uncaptioned image]satellite images only). While hydrometeorological data (setting [Uncaptioned image]environmental predictors only) are commonly used in global and continental wildfire forecasting models, they can be complemented by satellite data to improve strongly both the spatial resolution and the accuracy of the prediction (setting [Uncaptioned image]satellite and environmental data), as reported in Table 4, where the multi-modal approach leads to the best performances.
@@ 5.2 Importance of Negative Sampling for Training and Evaluation
Numerous wildfire forecasting benchmarks require sampling the negative (non-fire) samples due to extreme imbalance and computational constraints. A common strategy is to focus on samples that burned once across the period studied (Bakke et al., 2023; Prapas et al., 2022, 2023). In CanadaFireSat, we opt for a different strategy: we sample negative examples for each Canadian province uniformly across their yearly fire driver patterns (FWI values). By sampling the training, validation, and test sets in this way, we aim to train and evaluate our models on a subset representative of the conditions encountered in all of Canada. Nonetheless, as the yearly fire patterns vary, the distribution of negatives with respect to the FWI changes over the years, in turn affecting the performance of models. This motivated the creation of a second test set to understand the impact of sampling on the models’ performance (Test Hard), where ignition, a complex phenomenon difficult to model (Chen et al., 2021; Calef et al., 2008), differentiates positive from negative patches. Indeed human-induced ignitions, generally caused by infrastructures, agricultural practices, or ”recreational” activities, are typically hard to estimate with CanadaFireSat, as Sentinel-2 is the only source providing information on human presence, but only for a limited spatial context of $2.64\>\text{km}\times 2.64\>\text{km}$. Fine-tuning our multi-modal models on data such as Test Hard or enhancing our set of predictors with proxies of ignition probability (e.g, proximity to human settlements, or lightning probability) are relevant directions for improving our models towards accounting for ignition probability.
@@ 5.3 Modeling Wildfires in the Boreal Ecosystem
As initially stated in Section 1, one of our main motivations is the rise of wildfires in the boreal ecosystem and the risks this represents for its local communities. To cover the areas of interest and to evaluate the broader impact of wildfires on global climate, we created our benchmark CanadaFireSat so that it covers the entirety of Canada, including all its agricultural lands, urban areas, and other ecosystems such as the temperate forest in British Columbia. To study the behavior of trained models on the boreal ecosystem, it is possible to constrain the analysis on the main land cover classes of the boreal ecosystem (needleleaf forest and wetlands). With our benchmark models trained on CanadaFireSat, we observe an important difference of performance between those two land cover classes: with Multi-Modal CNN and ViT performing respectively $11.9\%$ and $11.4\%$ better on needleleaf forest compared to wetlands in terms of weighted F1 score on both the Val and Test sets, showing that for the latter land cover, performance is still not optimal. Indeed, wetland wildfires are a unique phenomenon compared to forest wildfires, as they depend much more on soil-related predictors and can burn underground for a long period. As a consequence, they are sometimes undetectable for optical remote sensing satellites. In particular, peatland wildfires that emit large amounts of CO2 and mercury (Fraser et al., 2018; Kohlenberg et al., 2018) are commonly studied independently from forest fires (Pelletier et al., 2023; Bali et al., 2021). Extending CanadaFireSat so that it includes data acquired from radar remote sensing satellites (for instance, Sentinel-1 images) could help to better model the surface soil conditions for wetlands (Millard and Richardson, 2018) and bridge the gap in performance across the boreal ecosystem.
@@ 5.4 Operationalization of the Model
Deploying models trained with CanadaFireSat over the entirety of Canada would require densely sampling the country with Sentinel-2 image time series, resulting in a huge amount of data to be processed. Indeed, the proposed dataset is aimed at modeling wildfire patterns at a moderate scale, but at high-resolution, and can be coupled with coarser resolution approaches (Prapas et al., 2022; Bali et al., 2021) to identify areas of interest and then apply our model to map such areas more precisely. Such coupling would allow wildfire management experts to target specific areas at risk for fine-grained wildfire forecasting or focus on areas that require more surveillance due to their proximity to local communities or due to their ecological and environmental interest. By alleviating the need for significant computational resources, it would break the barrier to scale this approach to large continental areas. Learning models directly capable of multi-scale prediction is an interesting future research direction to deploy high-resolution wildfire forecasting at scale. Such a model would exploit hierarchical learning approaches developed in computer vision for semantic segmentation (Li et al., 2022; Atigh et al., 2022).
@@ 5.5 Limitations and Future Work
As mentioned in Section 5.2, the main limitation of methods trained on CanadaFireSat is the difficulty of modeling the ignition component in wildfires due to its inherent stochasticity. Weather data from ERA5 can provide information on the risk of lightning, nonetheless, explicitly adding lightning probability (Geng et al., 2019) as a predictor, as well as other proxies for human ignition like the proximity to human settlement could help the trained models to better characterize ignition.
@@
Multi-task learning (Zhang and Yang, 2021) could also be leveraged to develop a model forecasting wildfires at multiple scales. One could leverage different forecasting heads at multiple resolutions: $10$ km, $1$ km, $100$ m. This could help alleviate memory size constraints when high-resolution forecasts are deemed unnecessary and help providing consistent predictions across scales.
@@
Moreover, one could investigate the potential of geolocation embeddings such as SatCLIP (Klemmer et al., 2023) or GeoCLIP (Vivanco Cepeda et al., 2023) to represent high-resolution non-dynamic satellite information. These could be combined with non-spatial, but temporal dynamics from Sentinel-2 (Pelletier et al., 2023) as a way to factorize spatial and temporal components in satellite data and limit memory consumption. Extending CanadaFireSat with atmospherically corrected images (e.g. L2A) or with BRDF-corrected Harmonized Landsat and Sentinel-2 data could help improving performances.
@@
Another line of future research deals with the improvement of the pretraining of our multi-modal deep learning approaches. In our work, we leverage image encoders pre-trained on natural images such as ImageNet or via DINOv2, which are very different from multi-spectral satellite images. With the drastic increase in availability of Earth observation data, several models are being proposed to learn in an unsupervised way generalizable representations from this data (Cong et al., 2022; Jakubik et al., 2023; Hong et al., 2024; Astruc et al., 2024; Sumbul et al., 2025). One could study the potential of those foundation models as pre-trained representations to be used in high-resolution wildfire forecasting; CanadaFireSat could be the perfect starting point for such an investigation.
@@
Finally, the increased complexity of models raises concerns regarding their interpretability and the possibility of understanding the role of the input variables in the final predictions. Several approaches exist to provide interpretations of black box wildfire forecasting models via feature attributions (Sundararajan et al., 2017; Selvaraju et al., 2017) or ranking (Lundberg and Lee, 2017), or even to directly build interpretable wildfire forecasting model architectures (Koh et al., 2020; Chen et al., 2019) via dense prediction architecture (Sacha et al., 2023; Porta et al., 2025a). However, those methods need adaptation to accommodate multi-modal (Ekim and Schmitt, 2023; Wang et al., 2023) or multi-temporal data (Turbé et al., 2023; Gee et al., 2019; Ghosal and Abbasi-Asl, 2021). They are also often not directly applicable to Earth observation data (Porta et al., 2025b) due to their strong implicit bias for natural images (Chen et al., 2019). This gap remains unfulfilled, and future works, for and beyond the wildfires prediction problem, should explore interpretable methods specifically tailored to Earth observation problems.